Miguel rented a truck for one day. There was a base fee of
$11.75
, and there was an additional charge of
7
cents for each mile driven. The total cost,
C
(in dollars), for driving
x
miles is given by the following function.
=Cx+0.07x11.75
What is the total rental cost if Miguel drove
40
miles?

Answer by Boreal(15235)

C(40)=11.75+40(0.07)=11.75+2.80=$14.55